OpenStudy (anonymous):

The circumference of a circle with a diameter of 3.73 in is _____in. Round answer to nearest hundredth

OpenStudy (campbell_st):

the formula is \[C = \pi \times d\] so you need to evaluate \[C = \pi \times 3.73\]

OpenStudy (anonymous):

i have the answer but i don't know how to round it

OpenStudy (anonymous):

could you show how to round it